OpenStudy (anonymous):

Still confused lol. Complete the following ordered pairs for the equation. 6x + -5y = 20 (-5, _) (_, -4) (_, 2)

OpenStudy (anonymous):

your points are ALWAYS in (x,y) (-5,_)....-5 = x now sub -5 in for x 6x + (-5y) = 20 6(-5) - 5y = 20 -30 - 5y = 20 -5y = 20 + 30 -5y = 50 y = -10 answer is (-5,-10) any questions ?

OpenStudy (anonymous):

Is it the signs that are confusing you ?

OpenStudy (anonymous):

Yes, like i see what you did and then i try it and get confused. so the second one i would start it off like 6(-4) +-5y=20 ? because i wasn't sure if i use +-5 or just do -5

OpenStudy (anonymous):

6x + (-5y) = 20 --- a + times a - = a negative. so you could just use this equation : 6x - 5y = 20 for the rest of your problems

OpenStudy (anonymous):

second one would start like this.. y = -4 6x - 5y = 20 6x - 5(-4) = 20 can you finish ? Don't forget a negative times a negative is a positive

OpenStudy (anonymous):

yeah i got 20 so would i minus that by 20? because that would just be zero and then my answer would just be zero? it's just so confusing